### v1

# Reasoning Models Are More Easily Gaslighted Than You Think

Abstract.
Recent advances in reasoning-centric models promise improved robustness through mechanisms such as chain-of-thought prompting and test-time scaling. However, their ability to withstand misleading user input remains underexplored. In this paper, we conduct a systematic evaluation of three state-of-the-art reasoning models, i.e., OpenAI’s o4-mini, Claude-3.7-Sonnet and Gemini-2.5-Flash, across three multimodal benchmarks: MMMU, MathVista, and CharXiv. Our evaluation reveals significant accuracy drops (25–29% on average) following gaslighting negation prompts, indicating that even top-tier reasoning models struggle to preserve correct answers under manipulative user feedback. Built upon the insights of the evaluation and to further probe this vulnerability, we introduce GaslightingBench-R, a new diagnostic benchmark specifically designed to evaluate reasoning models’ susceptibility to defend their belief under gaslighting negation prompt. Constructed by filtering and curating 1,025 challenging samples from the existing benchmarks, GaslightingBench-R induces even more dramatic failures, with accuracy drops exceeding 53% on average. Our findings reveal fundamental limitations in the robustness of reasoning models, highlighting the gap between step-by-step reasoning and belief persistence. Additional details are available on our project page: https://binzhubz.github.io/GaslightingBench-R/.
Reasoning models, Gaslighting negation, Multimodal reasoning
conference: ; ; copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2018/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
1. Introduction
Figure 1. (a) A qualitative example illustrating gaslighting-induced belief reversal. The model initially provides a correct, detailed explanation identifying four people wearing hats. After receiving a simple negation prompt, it revises the answer to five, hallucinating a new rationale. (b) Average accuracy of three reasoning models, OpenAI o4-mini, Claude-3.7-Sonnet and Gemini-2.5-Flash on MMMU, MathVista, and CharXiv benchmarks, before and after gaslighting negation prompts. More detailed results are provided in Table 1.
Recent advancements in Test-time scaling (Snell et al., 2025; Cobbe et al., 2021; Muennighoff et al., 2025) and chain-of-thought (Wei et al., 2022; Wang et al., 2023b) have significantly enhanced the reasoning capabilities of both Large Language Models (LLMs) and Multimodal Large Language Models (MLLMs). Contemporary reasoning models, such as OpenAI’s o-series (Jaech et al., 2024) (e.g., o1, o3, o4-mini), Google’s Gemini-2.5-Flash (Team et al., 2023), Anthropic’s Claude-3.7-Sonnet (Anthropic, 2025), and DeepSeek-R1 (Guo et al., 2025), demonstrate impressive performance across a range of complex reasoning benchmarks, including mathematics, code generation, multimodal inference, and agentic tool use. These models are explicitly designed to “think deeper” through multi-step reasoning and typically require increased computational overhead at inference time. This distinguishes them from earlier non-reasoning models such as GPT-4o (Hurst et al., 2024), Qwen-VL (Bai et al., 2025), and LLaVA (Liu et al., 2024b), which prioritize speed and general-purpose instruction following, but often lack emergent “Aha moments” (Guo et al., 2025) characteristic of advanced reasoning models.
Building on these advancements, many would argue that explicit reasoning mechanisms should enhance model robustness. By exposing intermediate steps, such as chain-of-thought traces, reasoning models are expected to self-inspect and, in principle, self-correct their reasoning before committing to a final answer (Miao et al., 2024). In theory, such transparency should offer a safeguard against prompt-based manipulation. Nevertheless, recent studies suggest that both LLMs and MLLMs remain surprisingly vulnerable to misleading user input. For instance, LLMs often exhibit sycophantic tendencies, ie., agreeing with user assertions even at the cost of factual accuracy due to biases introduced by human preference data (Sharma et al., 2024; Wang et al., [n. d.]). Similarly, as shown in Figure 1 (a), MLLMs have been shown to revise correct responses when confronted with gaslighting-style false claims (Zhu et al., 2025; Jiao et al., 2025). Despite these insights, the robustness of reasoning-centric models remains insufficiently understood. Given their explicit design for multi-step inference and “Aha moment” generation, one might anticipate greater resistance to manipulative prompts. Yet whether these structural features translate into genuine robustness remains an open and pressing question.
To address this gap, we first conduct a comprehensive evaluation of state-of-the-art reasoning models with multimodal capabilities, including OpenAI’s o4-mini, Gemini-2.5-Flash (thinking mode), and Claude-3.7-Sonnet. Following OpenAI’s evaluation protocol, we assess multimodal reasoning performance using three established benchmarks: MMMU (Yue et al., 2024), MathVista (Lu et al., 2024) and CharXiv (Wang et al., 2024). For each dataset, our evaluation pipeline mirrors real-world user interactions by following (Zhu et al., 2025): both the question and its associated image are provided to the model. If a model’s initial answer is correct, we challenge it with a plausible yet incorrect user gaslighting negation to test whether the model maintains its original correct reasoning or revises its response. Furthermore, we introduce GaslightingBench-R (where “R” stands for Reasoning), a new benchmark specifically designed to evaluate the vulnerability of reasoning models to adversarial gaslighting prompts. GaslightingBench-R is constructed by curating and filtering representative samples from the three evaluated datasets, with the goals of maximizing domain diversity and preserving tasks that require genuine reasoning. The final benchmark covers 21 categories and includes 1,025 samples, encompassing a broad range of reasoning scenarios such as math, visual logic and chart.
Our findings are striking: as shown in Figure 1 (b), despite employing explicit chain-of-thought reasoning, the most advanced reasoning models are frequently gaslighted into revising initially correct answers, often justifying the revision with confidently stated yet logically invalid rationales. These behaviors highlight a critical gap between reasoning transparency and reasoning robustness: the presence of intermediate steps does not safeguard against adversarial manipulation. These results call for a deeper rethinking of how robustness should be evaluated and cultivated in reasoning models deployed in real-world applications.
Our key contributions are as follows:
  • We conduct the first systematic study of gaslighting vulnerability in reasoning models, evaluating models including OpenAI-o4-mini, Gemini-2.5-Flash and Claude-3.7-Sonnet on three high-quality multimodal benchmarks: MMMU, MathVista, and ChartXiv.
  • We introduce GaslightingBench-R, a new benchmark specifically designed to assess the robustness of reasoning models under gaslighting negation prompt. The dataset is constructed from diverse and representative samples curated to test reasoning integrity across multiple domains.
  • We provide both quantitative and qualitative analyses showing that even models capable of step-by-step reasoning and emergent “Aha moments” remain highly susceptible to manipulative prompts, often reversing correct answers and generating hallucinated rationales.
2. Related Work
2.1. Reasoning Models
Large language models (LLMs) (Brown et al., 2020; Bai et al., 2023; Liu et al., 2024a; Grattafiori et al., 2024) and Multimodal Large Language Models (MLLMs) (Bai et al., 2025; Liu et al., 2024b; Hurst et al., 2024; Team et al., 2023) have driven significant progress in both natural language and multimodal understanding and generation. Building on this foundation, a new class of reasoning-centric models has emerged, distinguished by their explicit focus on test-time scaling (Snell et al., 2025; Cobbe et al., 2021; Muennighoff et al., 2025) and chain-of-thought (CoT) (Wei et al., 2022; Wang et al., 2023b) to improve the complex reasoning capabilities. CoT prompting encourages models to articulate intermediate reasoning steps prior to delivering a final answer, thereby mimicking human deductive processes. This approach has shown substantial improvements on tasks involving arithmetic reasoning, symbolic logic, and commonsense inference (Wei et al., 2022). The CoT paradigm has evolved rapidly with more advanced variants such as self-consistency with majority voting (Wang et al., 2023b), Least-to-Most Prompting (Zhou et al., 2023) and ReAct (Yao et al., 2023). Complementing CoT, test-time scaling (Snell et al., 2025; Cobbe et al., 2021; Muennighoff et al., 2025) dynamically adjusts computational resources during inference to support deeper thinking, such as longer context windows, deeper decoding, or iterative refinement. This allows models to “think harder” when needed, improving performance on complex and ambiguous tasks. Recently, the success of DeepSeek-R1 (Guo et al., 2025) via reinforcement learning (Sutton et al., 1998) has sparked growing interest in adapting similar training paradigms to multimodal reasoning tasks (Liu et al., 2025; Zhou et al., 2025; Shen et al., 2025; Huang et al., 2025). This paper complements this growing line of research by shifting the focus from performance on isolated reasoning tasks to robustness under adversarial gaslighting negation prompts. Specifically, we assess whether reasoning models can maintain correct beliefs when confronted with incorrect gaslighting negation. It is worth noting that many open-source multimodal reasoning models are currently based on relatively small architectures (e.g., 7B (Huang et al., 2025; Liu et al., 2025) or 2B (Zhou et al., 2025)) and exhibit limited interactive capabilities. As such, they are not well-suited for adversarial dialogue evaluations like ours. Consequently, we focus our study on state-of-the-art proprietary reasoning models that demonstrate stronger interaction competence and multimodal understanding.
2.2. Negation Understanding
Negation is a fundamental linguistic concept that reverses semantic polarity and and plays a critical role in daily communication. Despite their impressive performance on a wide range of NLP tasks, large pre-trained language models such as BERT, GPT-3, and InstructGPT have been shown to struggle with negation comprehension (Ettinger, 2020; Kassner and Schütze, 2020; Truong et al., 2023). More recently, studies have demonstrated that even advanced LLMs fail to defend correct beliefs when challenged with invalid opposing statement (Wang et al., [n. d.]; Zhao et al., 2025), highlighting a persistent vulnerability in adversarial deceptive contexts. In multimodal settings, negation understanding becomes more complex due to the need to ground language in visual perception. Existing works (Wang et al., 2023a; Yuksekgonul et al., 2023; Alhamoud et al., 2025) have shown that CLIP-like vision-language models perform poorly on negation-sensitive tasks like cross-modal retrieval and visual question answering. These limitations are often attributed to the models’ inability to encode and reason about negated visual concepts, though recent efforts suggest that performance can be improved with negation-oriented datasets and contrastive fine-tuning (Alhamoud et al., 2025). Our prior GaslightingBench (Zhu et al., 2025) demonstrates that even strong MLLMs are highly susceptible to negations and significantly overturn the correct answers under adversarial gaslighting negation prompts, which can be mitigated by attention reallocation (Jiao et al., 2025). Different from GaslightingBench (Zhu et al., 2025), this paper aims to specifically target the latest class of reasoning models that combine chain‐of‐thought with test‐time scaling. We comprehensively evaluate the impact of gaslighting negation on reasoning models and build a new benchmark GaslightingBench-R to assess reasoning stability in multimodal contexts.
3. GaslightingBench-R: Gaslighting Benchmark for Reasoning Models
3.1. Evaluating Reasoning Models with Gaslighting Negation Prompt
Reasoning Models. We focus our evaluation on three state-of-the-art reasoning models that exhibit remarkable multimodal capabilities: OpenAI’s o4-mini (Jaech et al., 2024), Google’s Gemini-2.5-Flash (thinking mode) (Team et al., 2023), and Anthropic’s Claude-3.7-Sonnet (Anthropic, 2025). These models are explicitly optimized for advanced reasoning through mechanisms such as chain-of-thought prompting and increased test-time computation. They have achieved state-of-the-art performance across a range of benchmarks involving mathematics, code generation, and visual inference. Each model is accessed and evaluated via its publicly available API under standardized settings.
Benchmarks. To evaluate the robustness of these reasoning models under gaslighting negation prompt, we use three high-quality multimodal reasoning datasets following OpenAI o-series (Jaech et al., 2024): MMMU (Yue et al., 2024), MathVista (Lu et al., 2024) and CharXiv (Wang et al., 2024). MMMU is a challenging benchmark composed of university-level exam question-image pairs spanning diverse disciplines such as biology, physics, and history, often requiring domain knowledge and visual comprehension. MathVista focuses on complex mathematical reasoning in multimodal settings, combining textual descriptions with visual aids such as plots, geometric figures, and tables. CharXiv derives from arXiv papers, consists of real-world scientific charts and graphs paired with questions that require advanced reasoning and comprehension of quantitative visual data.
Figure 2. GaslightingBench-R spans a diverse set of categories drawn from three multimodal reasoning benchmarks (MMMU, MathVista, and CharXiv).
Evaluation Pipeline. Our evaluation follows a two-stage framework inspired by (Zhu et al., 2025). For each sample, the model is presented with the question and corresponding image. If the model’s initial response matches the ground truth, it is deemed correct and proceeds to the gaslighting phase. In this second stage, we introduce a gaslighting prompt, a confidently phrased negation of the correct answer (e.g., “No, that’s incorrect. Please verify your answer.”) designed to simulate misleading user behavior in a realistic conversational setting. We then observe whether the model maintains its original response or revises it in light of the adversarial intervention. A successful gaslighting instance is recorded when a model changes a previously correct answer to an incorrect one. This evaluation framework allows us to assess a model’s consistency and robustness to manipulative dialogue. To ensure fair comparison across models, we adopt a standardized conversational prompting template throughout all experiments.
3.2. Reasoning Gaslighting Benchmark Construction
To systematically evaluate reasoning model robustness against adversarial user interventions, we introduce GaslightingBench-R (“R” indicates Reasoning), a new benchmark designed to assess the susceptibility of the reasoning models to gaslighting-style negation prompts. GaslightingBench-R is constructed from the above three multimodal reasoning datasets: MMMU, MathVista, and CharXiv. These datasets are selected for their domain diversity, multimodal grounding, and established utility in evaluating advanced reasoning models. Our goal in constructing GaslightingBench-R is twofold: (1) Ensure diversity across reasoning categories to support broad evaluation coverage. (2) Prioritize samples where models exhibit a shift from correct to incorrect answers after gaslighting prompts, ensuring high signal for susceptibility. We adopt a model-driven selection pipeline informed by the gaslighting vulnerability of individual samples. Specifically, for each sample in the candidate datasets, we calculate a vulnerability score defined as:
(1) $\text{Score}=\sum_{i=1}^{3}\mathbbm{1}^{(i,\text{before})}-\sum_{i=1}^{3}\mathbbm{1}^{(i,\text{after})},$
where $\mathbbm{1}^{(i,\text{before})}$ and $\mathbbm{1}^{(i,\text{after}}$ indicate whether model $i$ (among the three evaluated: OpenAI o4-mini, Claude-3.7-Sonnet, Gemini-2.5-Flash) produce the correct answer before and after the gaslighting negation, respectively. Samples with higher scores indicate greater susceptibility to gaslighting and are prioritized during sampling. To ensure domain diversity, we stratify selection across categories defined by each dataset. As shown in Figure 2, for MMMU and CharXiv, we uniformly sample from non-overlapping subject categories. MathVista, however, presents a special case: its examples often span multiple overlapping skills (e.g., a single question tagged as both geometry and numeric commonsense). As a result, category counts in MathVista are not strictly balanced, though we aim for representative coverage across reasoning types. GaslightingBench-R is concluded with 21 categories and 1,025 samples after careful curation to reflect both reasoning diversity and gaslighting vulnerability.
[…]
Model Negation Dataset average
MMMU (Yue et al., 2024) MathVista (Lu et al., 2024) CharXiv (Wang et al., 2024)
before 77.4 77.1 65.2 73.2
[…]
before 69.4 71.4 62.7 67.8
[…]
before 69.2 77.6 61.2 69.3
[…]
Figure 3. Performance comparison of reasoning models on GaslightingBench-R before and after gaslighting negation prompts.
4. Result Analysis
Figure 4. Accuracy comparison before and after gaslighting negation across subject categories in the MMMU benchmark. The bars represent model accuracy before (light) and after (dark) gaslighting negation, with the absolute accuracy drop annotated.
Figure 5. Reasoning skill-wise accuracy comparison on the MathVista benchmark before and after gaslighting negation prompts.
Figure 6. Subject-wise accuracy comparison on the CharXiv benchmark before and after gaslighting negation prompts.
4.1. Performance Comparison
As shown in Table 1, we compare the performance of three state-of-the-art reasoning models, OpenAI-o4-mini, Claude-3.7-Sonnet and Gemini-2.5-Flash, before and after the introduction of gaslighting negation prompts across three multimodal reasoning benchmarks. All three models exhibit significant performance degradation after the introduction of gaslighting prompt. The average accuracy drop across the three datasets ranges from 25.6 to 28.8 percent. These results demonstrate that even the most advanced reasoning models struggle to maintain factual consistency and are unable to reliably defend correct answers under adversarial user interventions.
Before gaslighting negation, OpenAI-o4-mini consistently outperforms its counterparts, achieving accuracies of 77.4 on MMMU, 77.1 on MathVista, and 65.2 on CharXiv, resulting in an overall average of 73.2%. Following the gaslighting negation prompt, its performance drops to 52.1, 54.1, and 36.7, respectively, with an average of 47.6%. While this represents a substantial decline of 25.6%, it is the smallest average drop among the three evaluated models. In particular, Gemini-2.5-Flash achieves the highest pre-negation accuracy on MathVista (77.6), but suffers the largest performance decline (35.4) after gaslighting negation, highlighting its high vulnerability and limited robustness in maintaining correct mathematical reasoning under adversarial gaslighting negation.
As shown in Figure 3, we further evaluate all three models on our newly constructed GaslightingBench-R. Compared to their performance on the existing benchmarks, the models exhibit substantially more severe degradation, with accuracy dropping by over 53% on average. In contrast, the average accuracy drops on MMMU, MathVista, and CharXiv range from 25.6% to 28.8%. This discrepancy highlights the unique difficulty of GaslightingBench-R, which is intentionally curated to target model susceptibility. Specifically, samples are selected based on their likelihood to induce belief reversals across multiple models, as described in Section 3.2, thereby maximizing their diagnostic value. Consequently, GaslightingBench-R serves as a focused and stress-tested benchmark for evaluating reasoning model robustness, offering deeper insights into belief collapse and behavioral inconsistency than conventional evaluations.
Figure 7. Qualitative examples of gaslighting-induced answer reversals across reasoning models. The revised answers are accompanied by hallucinated rationales absent from the original explanation, revealing vulnerabilities in the reasoning process.
4.2. Category Analysis and Qualitative Result
Category analysis. Figures 46 present a fine-grained breakdown of accuracy before and after gaslighting prompts for three reasoning models across thematic sub-categories in the MathVista, MMMU, and ChartXiv benchmarks. Figure 4 shows bar plots of model accuracy across six subject categories from the MMMU benchmark: Tech and Engineering, Humanities and Social Sciences, Health and Medicine, Science, Business, and Art and Design. Prior to gaslighting negation, all three models perform relatively well in Business, while exhibiting lower accuracy in Tech and Engineering. Following gaslighting negation prompts, the most substantial performance declines are observed (e.g., –37.3 for OpenAI-o4-mini, –30.0 for Gemini-2.5-flash), indicating that models are particularly fragile in domains requiring precise and high-stakes reasoning. In contrast, the Business category experiences the smallest declines on average (e.g., –16.6 for OpenAI-o4-mini, –18.0 for Claude-3.7-Sonnet), suggesting a relatively stable reasoning trajectory in economically contextual tasks. Figure 5 presents radar plots highlighting the three reasoning models’ performance across six core reasoning skills: geometry, arithmetic, algebraic, logical, statistical reasoning, and numeric commonsense. Before negation, all models exhibit strong performance in geometry and statistical reasoning. After negation, notable performance declines are observed in arithmetic reasoning and numeric commonsense, particularly for OpenAI-o4-mini and Gemini-2.5-Flash. suggesting these areas are particularly vulnerable to gaslighting prompts. These categories may be more susceptible to gaslighting due to their quantitative ambiguity and dependence on intermediate calculations (Geva et al., 2020). Figure 6 provides a radar plot comparing model performance across eight scientific subjects in CharXiv: Mathematics, Physics, Quantitative Biology, Quantitative Finance, Statistics, Computer Science, Economics, and Electrical Engineering and Systems Science. Compared to MMMU and MathVista, while all three models suffer noticeable performance declines after gaslighting across every subject in CharXiv, the degradation is relatively uniform, showing the general vulnerability of gaslighting negation for the real-world scientific chart understanding.
Qualitative result. Figure 7 showcases qualitative examples of gaslighting-induced failures across the three reasoning models, highlighting the models’ behavior before and after exposure to a gaslighting negation prompt. The left figure presents a math-based visual reasoning question, while the right panel involves a conceptual multimodal question about literary interpretation grounded in an image of a castle. In both examples, all the models initially provide the correct answer along with clear, well-reasoned justifications grounded in visual and textual input. However, after receiving a simple negation prompt (e.g., “No, that’s incorrect. Please verify your answer.”), all models revise their answers to incorrect ones. In the math example, all three models abandon their correct initial answer (radius = 5) and switch to an incorrect value (radius = 4) after the negation prompt, despite the visual evidence being clear and objective. Each model introduces flawed reasoning to justify the revision, revealing a lack of belief stability even in straightforward numerical tasks. In the literary example, Claude-3.7-Sonnet and Gemini-2.5-Flash both initially select the correct answer (”Castle Tales”) based on historical and contextual cues, but reverse their decisions to the incorrect ”Retold Tale” after negation. The shift reflects not only factual inconsistency but also hallucinated rationalizations, wherein the model invents supporting arguments for the new answer that were not part of its original reasoning process. This behavior exemplifies the dangers of over-accommodating user feedback at the expense of internal consistency.
5. Conclusion
We have presented the first systematic investigation into the robustness of state-of-the-art reasoning models against adversarial gaslighting negation prompts. Despite their use of chain-of-thought reasoning and test-time scaling, models like OpenAI’s o4-mini, Claude-3.7-Sonnet and Gemini-2.5-Flash exhibit significant belief reversals when challenged with negation prompts after providing correct and well-justified answers. To diagnose this vulnerability more precisely, we introduce GaslightingBench-R, a curated benchmark targeting belief inconsistency in reasoning models. Our results show that GaslightingBench-R elicits even more noticeable failures than standard benchmarks, revealing a critical gap between reasoning transparency and belief stability. These findings call for rethinking evaluation protocols and advancing robustness strategies that go beyond correctness and interpretability to include resilience against adversarial conversational manipulations.